Does the equation y/x^2 = 13.9 represent direct variation of one variable with the square of another variable? Write Yes or No.
Answer by venugopalramana(3286)

Does the equation y/x^2 = 13.9 represent direct variation of one variable with the square of another variable? Write Yes or No.
Y=13.9*X^2=K*X^2...WHERE K IS A CONSTANT..HENCE..
Y VARIES DIRECTLY AS SQUARE OF X ..SO ANSWER IS ..."YES"